OpenStudy (anonymous):

1/cot theta + cot theta = csc^2 thetacot theta

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{1}{cotθ} +cotθ=csc^2θ}\) are you solving or verifying ?

OpenStudy (anonymous):

VERIFYING I THINK

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{1}{cotθ} +cotθ=csc^2θ}\) \(\Huge\color{blue}{ \sf \frac{1}{cotθ} + \frac{cot^2θ}{cotθ}=csc^2θ}\) do you agree with this step?

OpenStudy (anonymous):

IDK I DONT GET IT'

OpenStudy (solomonzelman):

I wrote cot x as cot^2x / cotx

OpenStudy (anonymous):

OH OK GOT IT

OpenStudy (solomonzelman):

cot x = cot x / 1 now I am going to multiply top and bottom by cot x and I get, cot ^2x/cot x

OpenStudy (solomonzelman):

OK,

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{1}{cotθ} + \frac{cot^2θ}{cotθ}=csc^2θ}\) add the fractions on the left side \(\Huge\color{blue}{ \sf \frac{1+cot^2θ}{cotθ}=csc^2θ}\) and we know that \(\Huge\color{red}{ \sf 1+cot^2θ=csc^2θ}\)

OpenStudy (anonymous):

IS THE RED THE ANSWER?

OpenStudy (solomonzelman):

No red is an identity.

OpenStudy (solomonzelman):

So the top of the fraction (on the left side) simplifies to csc^2x so our next step to simplify the equation would be \(\Huge\color{blue}{ \sf \frac{csc^2θ}{cot^2θ}=csc^2θ }\) see how i got this ?

OpenStudy (anonymous):

do i write that for the answer?

OpenStudy (anonymous):

yes got it

OpenStudy (solomonzelman):

well if you are verifying you would say that it is NOT an identity. because \(\Huge\color{blue}{ \sf \frac{csc^2θ}{cot^2θ}≠csc^2θ }\)

OpenStudy (solomonzelman):

but if you are solving for x, you would divide both sides by csc^2θ \(\Huge\color{blue}{ \sf \frac{csc^2θ}{cot^2θ}=csc^2θ }\) \(\Huge\color{blue}{ \sf \frac{1}{cot^2θ}=1 }\) \(\Huge\color{blue}{ \sf tan^2θ=1 }\) and then you can take it from here to solve for x.

OpenStudy (solomonzelman):

That is if you need to solve. for verifying I already showed you.

OpenStudy (anonymous):

sooo what would be the answer for verifying?

OpenStudy (solomonzelman):

for verifying, NOT AN IDENTITY

OpenStudy (anonymous):

yes verifying

OpenStudy (anonymous):

i got more problems

OpenStudy (anonymous):

(1+ cos theta) (1-cos theta) = sin^2 theta